OpenStudy (anonymous):

how to determine a quadratic function if table of values is only the given??? for example (0,3)(1,4) (2,7) (3,12) (4,19) can you help me to find the quadratic function??? please!!!!

OpenStudy (imstuck):

When I graphed those points what I got was half of a parabola with the vertex at (0, 3).If you do not have any other points like the focus, or any other info like the directrix, we have to assume that you need just the standard form of the equation for a parabola centered at (0, 3) which is \[y-3=x ^{2}\]. That is the parabola with the vertex at (0, 3). That's the best I can do without a focus or something like that that is a bit more informative.

OpenStudy (anonymous):

can anyone give me the whole answer.. :3

OpenStudy (imstuck):

Do you have choices for the equation? Without the location of the focus or the line of directrix, there's no way you will find the whole equation.

OpenStudy (anonymous):

OUR TEACHER JUST GAVE THIS retrice TO US AND I DONT HAVE AN IDEA HOW TO DO IT.. BUT I KNOW HOW TO MAKE A TABLE FROM QUADRATIC FUNCTION, I JUST DONT KNOW HOW TO MAKE A QUADRATIC FUNCTION OUT OF TABLE OF VALUES..

OpenStudy (imstuck):

Ok, I got it, and the answer I gave you above it the correct one. The equation is x^2 = (y - 3). The focus is at (0, 3.25) and the line of directrix is y = 2.75. That is the right answer.